OpenStudy (yacoub1993):

Graph the function. y = - 3/4 sin (x - pi/4)

OpenStudy (yacoub1993):

\[y=-\frac{ 3 }{ 4 } \sin(x-\frac{ \pi }{ 4 })\]

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} &\textit{amplitude}&\textit{horizontal shift}\\ &\quad \downarrow&\qquad\qquad\downarrow \\ y=&-\cfrac{3}{4}sin&\left(x\qquad-\cfrac{\pi}{4}\right)\\ \end{array}\)

OpenStudy (jdoe0001):

use the graph of the parent sin(x) and "shift it" accordingly

OpenStudy (anonymous):

sure

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=ijTIr-aykUk

OpenStudy (anonymous):

f1q1g5 is the one you want. I can tell because your modified sin function has a zero at pi/4, and the only answer you posted to have that is f1q1g5